OpenStudy (anonymous):

Eighty ounces of Lazy Lawn fertilizer covers 1,250 squares feet of lawn. a. How many ounces would be required to cover a 4,000 square foot lawn?

OpenStudy (anonymous):

you can set this up as a proportion 8/1250 = x/4000 ---8 oz to 1250 ft = x oz to 4000 ft cross multiply 1250x = (4000)(8) 1250x = 32000 --- divide by 1250 x = 32000/1250 x = 25.6 25.6 oz is needed for 4000 ft